# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Bulk-Surface Interaction

### v2

@@ Abstract
We derive a thermodynamically consistent model, which describes the time evolution of a two-phase flow in an evolving domain. The movement of the free boundary of the domain is driven by the velocity field of the mixture in the bulk, which is determined by a Navier–Stokes equation. In order to take interactions between bulk and boundary into account, we further consider two materials on the boundary, which may be the same or different materials as those in the bulk. The bulk and the surface materials are represented by respective phase-fields, whose time evolution is described by a bulk-surface convective Cahn–Hilliard equation. This approach allows for a transfer of material between bulk and surface as well as variable contact angles between the diffuse interface in the bulk and the boundary of the domain. To provide a more accurate description of the corresponding contact line motion, we include a generalized Navier slip boundary condition on the velocity field. We derive our model from scratch by considering local mass balance and energy dissipation laws. Finally, the derivation is completed via the Langrange multiplier approach. We further show that our model generalizes previous models from the literature, which can be recovered from our system by either dropping the dynamic boundary conditions or assuming a static boundary of the domain.
@@ 1  Introduction
The description of two-phase flows is an important topic in modern continuum fluid dynamics, which has widespread applications in biology, chemistry, and engineering. To represent the interface separating the different components, two fundamental approaches have been developed, namely sharp-interface methods and diffuse interface methods. We refer to [32, 3, 46, 72] for a comparison of the two approaches.
@@
In diffuse interface models, the interface between the two materials is represented by a thin layer whose thickness is proportional to a small parameter $\varepsilon>0$. The location of the two components is usually represented by an order parameter, the so-called phase-field, which describes the difference of the local concentrations (or volume fractions) of the materials. This means that the phase-field attains values close to 1-1 or $1$ in the regions where the single fluids are present. At the the diffuse interface between the fluids, the phase-field is expected to exhibit a continuous transition between the values representing the pure phases. In normal direction across the diffuse interface, a certain transition profile can usually be observed. The main advantage of this approach is that the time evolution of the phase-field can be described via an Eulerian formulation as a PDE on the whole domain. This avoids directly tracking the interface between the phases. If a proper scaling with respect to the interface parameter $\varepsilon$ is used, the corresponding sharp-interface model can be recovered (in most cases at least formally) as the sharp-interface limit (see, e.g., [3]).
@@
One of the most fundamental models for describing the motion of two viscous and incompressible fluids with equal densities is the Model H. It was first proposed in [56] and later rigorously derived in [53]. The system consists of an incompressible Navier–Stokes equation that is coupled to a convective Cahn–Hilliard equation. The fundamental assumption in the derivation of the Model H is that the individual constant densities of the fluids are equal. They are therefore referred to as matched densities. This directly entails that the density of the whole mixture is constant.
@@
In the past two decades, various works have proposed generalizations of the Model H describing incompressible mixtures with unmatched densities and/or compressible two-phase flows. We refer to [5, 65, 19, 31, 75, 76, 46, 55, 37] to at least mention some of them.
@@
One of the most popular models for viscous incompressible two-phase flows with unmatched densities is the one derived by Abels, Garcke and Grün in [5]. In view of the initials of the authors, it is often referred to as the AGG Model. It was shown in [5] that by including an additional flux term in the Navier–Stokes equation, the model can be made thermodynamically consistent even if the constant individual densities of the fluids are unmatched. Analytical results for the AGG model can be found, e.g., in [1, 2, 9, 4, 6, 48, 49]. An alternative derivation of the AGG model via mixture theory was recently developed in [77].
@@
However, as the classical AGG model is considered subject to a no-slip boundary condition for the velocity field as well as homogeneous Neumann boundary conditions for the Cahn–Hilliard quantities (i.e., the phase-field and the chemical potential), it still has some drawbacks if a precise description of the behavior of the fluids close to the boundary is required. In the Cahn–Hilliard equation, the homogeneous Neumann condition on the phase-field enforces the diffuse interface to always intersect the boundary at a ninety degree contact angle, which is very unrealistic in many applications. The homogeneous Neumann boundary condition on the chemical potential entails the conservation of mass within the considered domain, which makes a description of transfer of material between bulk an boundary (e.g., absorption processes) impossible. These issues have also been noticed by physicists (see, e.g., [18, 35, 36, 57]), who proposed the inclusion of an additional surface free energy to describe short-range interactions between bulk and boundary more precisely.
@@
Due to these observations, several types of dynamic boundary conditions for the Cahn–Hilliard equation have been proposed and analyzed in the literature. Especially in recent times, dynamic boundary conditions that also exhibit a Cahn–Hilliard type structure have become very popular because they allow for a quantification of the mass transfer between bulk and boundary. For brevity, we refer to such models as bulk-surface Cahn–Hilliard models. We refer to [43, 44, 27, 26, 25, 38, 71] for analytical results and to [68, 69, 54, 67, 14, 15, 21, 20] for a numerical investigation of the bulk-surface Cahn–Hilliard system. A nonlocal bulk-surface Cahn–Hilliard system was analyzed in [60, 66]. A convective variant of the bulk-surface Cahn–Hilliard system with a prescribed velocity field was analyzed, for example, in [28, 29, 47, 51, 61, 62]. The role of dynamic boundary conditions in Cahn–Hilliard type models will be discussed in more detail in Subsection 2.3. For more information on the Cahn–Hilliard equation with classical homogeneous Neumann boundary conditions or with dynamic boundary conditions, we refer to the recent review paper [78] as well as the book [70].
@@
Moreover, as shown for example in [73, 74], a no-slip boundary condition on the velocity field is not very well suited for the description of moving contact line phenomena. This is because convective effects occurring at the boundary (or close to the boundary) are neglected if the velocity field is assumed to be identically zero at the surface. Therefore, if a no-slip boundary condition is imposed, the motion of the contact line is primarily driven by diffusion or convection processes away from the boundary. To overcome this issue, a generalized Navier slip boundary condition, which provides a better description of moving contact lines, was derived in [73].
@@
Based on these considerations, Navier–Stokes–Cahn–Hilliard models with dynamic boundary conditions on the Cahn–Hilliard subsystem as well as a generalized Navier slip boundary condition for the Navier–Stokes equation have been proposed and investigated in the literature. We refer to [40, 41], where a second-order (Allen-Cahn type) dynamic boundary condition was used for the Cahn–Hilliard subsystem. Here, the Allen-Cahn type dynamic boundary condition can be interpreted as a parabolic relaxation of a transport-type dynamic boundary condition related to contact angle dynamics (see, e.g., [80]). Instead, in [50], a fourth-order (Cahn–Hilliard type) dynamic boundary condition is considered, which allows for a better description of the transfer of material between bulk and boundary. This model was further analyzed in [42]. In these models, the Cahn–Hilliard quantities at the boundary enter the generalized Navier slip boundary condition via a forcing term. We also refer to [34, 24], where compressible two-phase flows were considered. In [34] a compressible Navier–Stokes–Allen–Cahn model was analyzed and in [24], a compressible Navier–Stokes–Cahn–Hilliard model with dynamic boundary conditions was investigated. Moreover, models for two-phase flows with phase-transition were studied in [10, 8].
@@
We point out that all the models discussed so far are each considered in a fixed bounded domain that does not change during the time evolution. In some applications, however, the domain in which a two-phase flow is considered will evolve over the course of time. A simple example is a droplet (surrounded by a further substance such as air) that is placed on a deformable substrate, for example, an elastic sheet (see Figure 1).
Figure 1: Droplet on a moving surface.
@@
Figure 2: Evolving domain (e.g., describing a bacterium) with
@@
In this paper, for the aforementioned reasons, we derive a new model, which generalizes the two-phase flow model derived in [50] to the framework of an evolving domain with a free boundary. Here, a key assumption is that the evolution of our domain is driven by the velocity field of the mixture contained in the domain. This means that the normal velocity of the free boundary is equal to the normal component of the velocity field at the boundary.
@@ 1.1  Formulation of our model
We further assume that the evolving domain $Q_{T}$ is transported by a vector field ${\mathbf{u}}\in C^{2}(\overline{Q_{T}})$, which is supposed to be the volume-averaged velocity field of the two materials that are contained inside the domain. This entails that $S_{T}$ is transported by the trace vector field ${\mathbf{u}}|_{S_{T}}\in C^{2}(S_{T})$. Since $S_{T}$ is orientable, there exists an evolving unit normal vector field ${\mathbf{n}}\in C^{2}(S_{T})$ and the normal velocity $\mathcal{V}$ of $S_{T}$ is thus given by
@@
For a mathematically precise formulation of the notion of evolving domains and surfaces that are transported by a velocity field, we refer to [13, 17].
@@
This system can be understood as a generalized version of the model introduced in [50], adapted to the situation of an evolving domain with free boundary. An equivalent reformulation of this model will be discussed in Subsection 2.1.
@@
The time evolution of $\mathbf{u}$} is described by the incompressible Navier–Stokes equations (1.1a). Here, $\rho(\phi)$ stands for the phase-field dependent density, $\mathbf{J}$ represents the mass flux related to interfacial motion, and $\mathbf{T}$ is the stress tensor. They are given by the formulas
@@
where the constants $\tilde{\rho}_{1},\tilde{\rho}_{2}>0$ are the individual densities of the fluids, $\nu(\phi)$ is the phase-field dependent viscosity, and
@@
denotes the symmetric gradient of $\mathbf{u}$}. The matrix $\mathbf{S}$ is referred to as the viscous stress tensor.
@@
to denote the material derivatives of $f$ and $g$, respectively. Here, $\tilde{g}$ denotes an arbitrary (sufficiently regular) extension of $g$ onto a neighborhood of $S_{T}$. In fact, the expression $\partial_{t}^{\circ}g$ does actually not depend on the choice of this extension (see, e.g., [13, Remark 2.7.2]). The phase-field dependent functions $m_{\Omega}(\phi)$ and $m_{\Gamma}(\psi)$ represent the mobilities of the bulk- and the surface materials, respectively. Moreover, $F^{\prime}$ and $G^{\prime}$ are the derivatives of double-well potentials $F$ and $G$, respectively. The bulk and surface quantities of the Cahn–Hilliard equations are further coupled through the boundary conditions in (1.1f), which are to be understood as follows:
@@
Here, $\alpha,\beta\in\mathbb{R}$ are prescribed parameters. The parameters $K,L\in[0,\infty]$ are used to distinguish different cases, each resulting in a certain solution behaviour related to a corresponding physical phenomenon (see Subsection 2.3 for more details). If $K=0$, we additionally assume that $\alpha\neq 0$.
Furthermore, on the surface $S_{T}$, the tangential component of the velocity field is governed by the generalized Navier slip boundary condition (1.1g) with a slip parameter $\gamma_{\tau}$ that may depend on the materials involved (i.e., the phase-fields). Here, for any vector field $\mathbf{w}$, the notation $\mathbf{w}_{\tau}$ indicates that only the tangential component of $\mathbf{w}$ is considered. The normal component of $\mathbf{u}$} is determined by the relation (1.1h), which involves the mean curvature $H=-\frac{1}{d}\textnormal{div}_{\Gamma}{\mathbf{n}}$. Moreover, $\gamma_{\mathbf{n}}$ is a prescribed scalar function that may depend on the phase-fields.
@@
    by means of the Reynolds transport theorem (see, e.g., [13, Theorem 2.11.1]). Similarly, as we also demand $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$, the transport theorem for evolving hypersurfaces (see, e.g., [13, Theorem 2.10.1]) yields
    $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{Vol}_{d-1}\big(\Gamma(t)\big)=\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Gamma(t)}1\;\mathrm{d}\mathcal{H}^{d-1}=\int_{\Gamma(t)}\textnormal{div}_{\Gamma}({\mathbf{u}})\;\mathrm{d}\mathcal{H}^{d-1}=0\quad\text{for all $t\in[0,T]$.}$ (1.14)
@@ 1.2  Further related literature
In contrast to our model, where the free boundary is the boundary of the evolving domain, there are also models for two-phase flows, where a free boundary contained in the interior of a static (i.e., time-independent) domain is considered. We refer, for example, to [45, 16]. A related sharp interface model with moving contact lines on inextensible elastic sheets was derived in [79]. For the investigation of Cahn–Hilliard type models on evolving surfaces without any bulk contribution, we refer to [7, 11, 12, 22, 23, 33] as well as the references therein.

### v3

@@ Abstract
We derive a thermodynamically consistent model, which describes the time evolution of a two-phase flow in an evolving domain. The movement of the free boundary of the domain is driven by the velocity field of the mixture in the bulk, which is determined by a Navier–Stokes equation. In order to take interactions between bulk and boundary into account, we further consider two materials on the boundary, which may be the same or different materials as those in the bulk. The bulk and the surface materials are represented by respective phase-fields, whose time evolution is described by a bulk-surface convective Cahn–Hilliard equation. This approach allows for a transfer of material between bulk and surface as well as variable contact angles between the diffuse interface in the bulk and the boundary of the domain. To provide a more accurate description of the corresponding contact line motion, we include a generalized Navier slip boundary condition on the velocity field. Based on local mass balance laws, we derive our model from scratch in two different ways: by the Lagrange Multiplier Approach and (in the case of matched densities and no mass flux between bulk and surface) by the Energetic Variational Approach. We further show that our model generalizes previous models from the literature, which can be recovered from our system by either dropping the dynamic boundary conditions or assuming a static boundary of the domain.
@@ 1  Introduction
The description of two-phase flows is an important topic in modern continuum fluid dynamics, which has widespread applications in biology, chemistry, and engineering. To represent the interface separating the different components, two fundamental approaches have been developed, namely sharp-interface methods and diffuse interface methods. We refer to [Du2020, Abels2018, Giga2018, PS2016] for a comparison of the two approaches.
@@
In diffuse interface models, the interface between two materials is represented by a thin layer whose thickness is proportional to a small parameter $\varepsilon>0$. The location of the two components is usually represented by an order parameter, the so-called phase-field, which describes the difference of the local concentrations (or volume fractions) of the materials. This means that the phase-field attains values close to 1-1 or $1$ in the regions where the single fluids are present. At the diffuse interface between fluids, the phase-field is expected to exhibit a continuous transition between the values representing the pure phases. In normal direction across the diffuse interface, a certain transition profile can usually be observed. The main advantage of this approach is that the time evolution of the phase-field can be described via an Eulerian formulation as a PDE on the whole domain. This avoids directly tracking the interface between the phases. If a proper scaling with respect to the interface parameter $\varepsilon$ is used, the corresponding sharp-interface model can be recovered (in most cases at least formally) as the sharp-interface limit (see, e.g., [Abels2018]).
@@
One of the most fundamental models for describing the motion of two viscous and incompressible fluids with equal densities is the Model H. It was first proposed in [Hohenberg1977] and later rigorously derived in [Gurtin1996]. The system consists of an incompressible Navier–Stokes equation that is coupled to a convective Cahn–Hilliard equation. The fundamental assumption in the derivation of the Model H is that the individual constant densities of fluids are equal. They are therefore referred to as matched densities. This directly entails that the density of the whole mixture is constant.
@@
In the past two decades, various works have proposed generalizations of the Model H describing incompressible mixtures with unmatched densities and/or compressible two-phase flows. We refer to [Abels2012, Lowengrub1998, Boyer2002, Ding2007, Shen2013, Shokrpour2018, Giga2018, Heida2012, Freistuhler2017] to at least mention some of them.
@@
One of the most popular models for viscous incompressible two-phase flows with unmatched densities is the one derived by Abels, Garcke and Grün in [Abels2012]. In view of the initials of the authors, it is often referred to as the AGG Model. It was shown in [Abels2012] that by including an additional flux term in the Navier–Stokes equation, the model can be made thermodynamically consistent even if the constant individual densities of fluids are unmatched. Analytical results for the AGG model can be found, e.g., in [Abels2013, Abels2013a, Abels2021, Abels2024, Abels2024a, Giorgini2021, Giorgini2022]. An alternative derivation of the AGG model via mixture theory was recently developed in [tenEikelder2023].
@@
However, as the classical AGG model is considered subject to a no-slip boundary condition for the velocity field as well as homogeneous Neumann boundary conditions for the Cahn–Hilliard quantities (i.e., the phase-field and the chemical potential), it still has some drawbacks if a precise description of the behavior of fluids close to the boundary is required. In the Cahn–Hilliard equation, the homogeneous Neumann condition on the phase-field enforces the diffuse interface to always intersect the boundary at a ninety degree contact angle, which is very unrealistic in many applications. The homogeneous Neumann boundary condition on the chemical potential entails the conservation of mass within the considered domain, which makes a description of transfer of material between bulk an boundary (e.g., absorption processes) impossible. These issues have also been noticed by physicists (see, e.g., [Binder1991, Fischer1997, Fischer1998, Kenzler2001]), who proposed the inclusion of an additional surface free energy to describe short-range interactions between bulk and boundary more precisely.
@@
Due to these observations, several types of dynamic boundary conditions for the Cahn–Hilliard equation have been proposed and analyzed in the literature. Especially in recent times, dynamic boundary conditions that also exhibit a Cahn–Hilliard type structure have become very popular because they allow for a quantification of the mass transfer between bulk and boundary. For brevity, we refer to such models as bulk-surface Cahn–Hilliard models. We refer to [Garcke2020, Garcke2022, Colli2020, Colli2022, Colli2022a, Fukao2021, Miranville2020] for analytical results and to [Metzger2021, Metzger2023, Harder2022, Meng2023, Bao2021, Bao2021a, Bullerjahn2024, Bullerjahn2025] for a numerical investigation of the bulk-surface Cahn–Hilliard system. A nonlocal bulk-surface Cahn–Hilliard system was analyzed in [Knopf2021b, Lv2025]. A convective variant of the bulk-surface Cahn–Hilliard system with a prescribed velocity field was analyzed, for example, in [Colli2018, Colli2019, Gilardi2019, Giorgini2025, Knopf2024, Knopf2024a]. The role of dynamic boundary conditions in Cahn–Hilliard type models will be discussed in more detail in Subsection 2.3. For more information on the Cahn–Hilliard equation with classical homogeneous Neumann boundary conditions or with dynamic boundary conditions, we refer to the recent review paper [Wu2022] as well as the book [Miranville-Book].
@@
Moreover, as shown for example in [Qian2006, Seppecher1996], a no-slip boundary condition on the velocity field is not very well suited for the description of moving contact line phenomena. This is because convective effects occurring at the boundary (or close to the boundary) are neglected if the velocity field is assumed to be identically zero at the surface. Therefore, if a no-slip boundary condition is imposed, the motion of the contact line is primarily driven by diffusion or convection processes away from the boundary. To overcome this issue, a generalized Navier slip boundary condition, which provides a better description of moving contact lines, was derived in [Qian2006].
@@
Based on these considerations, Navier–Stokes–Cahn–Hilliard models with dynamic boundary conditions on the Cahn–Hilliard subsystem as well as a generalized Navier slip boundary condition for the Navier–Stokes equation have been proposed and investigated in the literature. We refer to [Gal2016, Gal2019], where a second-order (Allen–Cahn type) dynamic boundary condition was used for the Cahn–Hilliard subsystem. Here, the Allen–Cahn type dynamic boundary condition can be interpreted as a parabolic relaxation of a transport-type dynamic boundary condition related to contact angle dynamics (see, e.g., [Yue2020]). Instead, in [Giorgini2023], a fourth-order (Cahn–Hilliard type) dynamic boundary condition is considered, which allows for a better description of the transfer of material between bulk and boundary. This model was further analyzed in [Gal2024]. In these models, the Cahn–Hilliard quantities at the boundary enter the generalized Navier slip boundary condition via a forcing term. We also refer to [Feireisl2010, Cherfils2019], where compressible two-phase flows were considered. In [Feireisl2010] a compressible Navier–Stokes–Allen–Cahn model was analyzed and in [Cherfils2019], a compressible Navier–Stokes–Cahn–Hilliard model with dynamic boundary conditions was investigated. Moreover, models for two-phase flows with phase-transition were studied in [Aki2014, Abels2025a].
@@
We point out that all the models discussed so far are each considered in a fixed bounded domain that does not change during the time evolution. In some applications, however, the domain in which a two-phase flow is considered will evolve over the course of time. A simple example is a droplet (surrounded by a further substance such as air) that is placed on a deformable substrate, for example, a deformable sheet (see Figure 1).
Abbildung 1: Droplet on a moving surface.
@@
Abbildung 2: Evolving domain (e.g., describing a bacterium) with
@@
In this paper, for the aforementioned reasons, we derive a new model, which generalizes the two-phase flow model derived in [Giorgini2023] to the framework of an evolving domain with a free boundary. Here, a key assumption is that the evolution of our domain is driven by the velocity field of the mixture contained in the domain. This means that the normal velocity of the free boundary is equal to the normal component of the velocity field at the boundary.
@@ 1.1  Formulation of our model
We further assume that the evolving domain $Q_{T}$ is transported by a sufficiently regular vector field ${\mathbf{u}}:\overline{Q_{T}}\to\mathbb{R}^{d}$, which is supposed to be the volume-averaged velocity field of the two materials that are contained inside the domain. This entails that $S_{T}$ is transported by the trace vector field ${\mathbf{u}}|_{S_{T}}:S_{T}\to\mathbb{R}^{d}$. Since $S_{T}$ is orientable, there exists an evolving unit normal vector field $\mathbf{n}$} and the normal velocity $\mathcal{V}$ of $S_{T}$ is thus given by
@@
For a mathematically precise formulation of the notion of evolving domains and surfaces that are transported by a velocity field, we refer to [BDGP-book, BGN-book].
@@
This system can be understood as a generalized version of the model introduced in [Giorgini2023], adapted to the situation of an evolving domain with free boundary. An equivalent reformulation of this model will be discussed in Subsection 2.1.
@@
The time evolution of $\mathbf{u}$} is described by the incompressible Navier–Stokes equations (1.1a). Here, $\rho(\phi)$ stands for the phase-field dependent density, $\mathbf{J}$ represents the mass flux related to interfacial motion, and $\mathbf{T}$ is the stress tensor. They are given by the formulae
@@
where the constants $\tilde{\rho}_{1},\tilde{\rho}_{2}>0$ are the individual densities of fluids, $\nu(\phi)$ is the phase-field dependent viscosity, and
@@
[…]
@@
to denote the material derivatives of $f$ and $g$, respectively. Here, $\tilde{g}$ denotes an arbitrary (sufficiently regular) extension of $g$ onto a neighborhood of $S_{T}$. In fact, the expression $\partial_{t}^{\circ}g$ does actually not depend on the choice of this extension (see, e.g., [BDGP-book, Remark 2.7.2]). The phase-field dependent functions $m_{\Omega}(\phi)$ and $m_{\Gamma}(\psi)$ represent the mobilities of the bulk- and the surface materials, respectively. Moreover, $F^{\prime}$ and $G^{\prime}$ are the derivatives of double-well potentials $F$ and $G$, respectively. The bulk and surface quantities of the Cahn–Hilliard equations are further coupled through the boundary conditions in (1.1f), which are to be understood as follows:
@@
Here, $\alpha,\beta\in\mathbb{R}$ are prescribed parameters. The parameters $K,L\in[0,\infty]$ are used to distinguish different cases, each resulting in a certain solution behaviour related to a corresponding physical phenomenon. They will be explained in more detail in Subsection 2.3. If $K=0$, we additionally assume that $\alpha\neq 0$.
[…]
@@
    by means of the Reynolds transport theorem (see, e.g., [BDGP-book, Theorem 2.11.1]). Similarly, as we also demand $\textnormal{div}_{\Gamma}{\mathbf{u}}=0$ on $S_{T}$, the transport theorem for evolving hypersurfaces (see, e.g., [BDGP-book, Theorem 2.10.1]) yields
    $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{Vol}_{d-1}\big(\Gamma(t)\big)=\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Gamma(t)}1\;\mathrm{d}\mathcal{H}^{d-1}=\int_{\Gamma(t)}\textnormal{div}_{\Gamma}({\mathbf{u}})\;\mathrm{d}\mathcal{H}^{d-1}=0\quad\text{for all $t\in[0,T]$}$ (1.14)
[…]
@@ 1.2  Further related literature
In contrast to our model, where the free boundary is the boundary of the evolving domain, there are also models for two-phase flows, where a free boundary contained in the interior of a static (i.e., time-independent) domain is considered. We refer, for example, to [Garcke2014, Barrett2017]. A related sharp interface model with moving contact lines on inextensible elastic sheets was derived in [YZR2023SharpInterface]. For the investigation of Cahn–Hilliard type models on evolving surfaces without any bulk contribution, we refer to [Abels2025, Bachini2023, Bachini2023a, Caetano2021, Caetano2023, Elliott2024] as well as the references therein.